OpenStudy (anonymous):

A steel sphere is dropped vertically onto a horizontal metal plate. The sphere hits the plate with a speed u, leaves it at speed v, and rebounds vertically to half its original height. Which expression gives the value of v/u?

OpenStudy (anonymous):

a) \[1/\sqrt{2} b) 1-1/\sqrt{2}\] c) 1/2^2 d) 1/2

OpenStudy (anonymous):

compare KE and PE from stage1 to KE and PE of stage 2 \[ \frac{{1\over2}mu^2}{{1\over2}mv^2}=\frac{mgh}{mg{h\over2}} \]

OpenStudy (anonymous):

Okay, so it v/u is 1/sqrt{2}?

OpenStudy (anonymous):

s